# Look-ahead Search on Top of Policy Networks in Imperfect Information Games

### 23 Dec 2023 v1

@@
Abstract.
Conducting additional search during test time is often used to improve the performance of reinforcement learning algorithms. Performing search in adversarial games with imperfect information is notoriously difficult and often requires a complicated training process. We present an algorithm that uses an arbitrary policy-gradient algorithm that learns from sampled trajectories in the setting of fully adversarial two-player games with imperfect information. Alongside the training of the policy network, the algorithm trains an additional critic network, which provides multiple expected values if both players follow one of a fixed set of transformations of the policy given by the policy network. These values are then used for depth-limited search. We show how the values from this critic can create a value function for imperfect information games. Moreover, they can be used to compute the summary statistics necessary to start the search from an arbitrary decision point in the game. The presented algorithm is scalable to very large games since it does not require any search in the training time. Furthermore, given sufficient computational resources, our algorithm may choose whether to use search or play the strategy according to the trained policy network anywhere in the game. We evaluate the algorithm’s performance when trained alongside Regularized Nash Dynamics, and we compare the performance of using the search against the policy network in the standard benchmark game of Leduc hold’em, multiple variants of imperfect information Goofspiel, and in a game of Battleships.
Key words and phrases:
Imperfect Information Games, Safe Search, Multi-agent Reinforcement Learning, Value Function, Multi-Valued States, Regularized Nash Dynamics, Counterfactual Regret Minimization, Multi-agent Search with Policy Transformation Value Network
1. Introduction
@@
Despite their success, the neural networks can occasionally lead the agent to make suboptimal decisions that an adversarial opponent may exploit. Wang et al.Wang et al. (2023) have illustrated how policy networks trained via AlphaZero-style algorithms in Go are susceptible to exploitation, as evidenced by winning against the trained policy network 97% of the time. Similarly, the authors of Perolat et al. (2022) report that their trained agents display evident gameplay errors, necessitating the application of multiple heuristics to rectify these errors. However, there are principled ways of using search to avoid these mistakes, as evidenced by AlphaZero Silver et al. (2018), DeepStack Moravčík et al. (2017), and Student of Games Schmid et al. (2023).
@@
Algorithms that use policy-gradient methods in imperfect information games like Regularized Nash Dynamics (RNaD) Perolat et al. (2022); Perolat et al. (2021) rely on directly applying the neural network’s policy during gameplay. Adding search to these methods could improve the strategy and mitigate these mistakes. It is important to note that this enhancement comes at the expense of transitioning previously model-free algorithms into a model-dependent during gameplay.
@@
Adding search to these policy-gradient methods in imperfect information games presents multiple challenges that are non-existent in a perfect information case. Firstly, the imperfect information search must be performed not only from the current state of the game but from all possible states that share the same public information amongst all players. Secondly, the state values are dependent on the strategies employed by each player. Thirdly, the opponent’s strategy is unknown, and the optimal strategies are not unique. This means that we need to optimize over all possible strategies the opponent may deploy both in the remainder of the game as well as unobserved past actions.
@@
In this paper, we introduce the algorithm Multi-agent Search with Policy Transformation Value Network (SePoT), which builds upon the concept of multi-valued states Brown et al. (2018). SePoT trains a policy deviation critic network, that estimates the expected value of policies derived from the policy network. This critic is trained alongside a policy network with a policy-gradient technique without requiring any additional trajectories beyond those sampled by the original policy-gradient algorithm. The critic network is trained with an off-policy V-trace that estimates expected values against a multitude of potential opponent strategies derived from the policy represented by the policy network. We show that this critic provides sufficient information to perform a safe search addressing all complications of search in imperfect information games.
@@
Furthermore, we present an approach that enables safe search at any point within a game, utilizing only the information provided by the introduced critic network and the policy network. This eliminates the need for conducting search in all the game parts from the root until the current decision point, a departure from prior methods Brown et al. (2018); Moravčík et al. (2017); Schmid et al. (2023). This flexibility allows us to perform search only in those game segments where we expect the strategy to be undertrained and we have enough computational resources while employing the neural network policy in other parts of the game.
@@
Finally, we establish a connection between value functions and multi-valued states within imperfect information games. With this connection, it becomes possible to train previously used imperfect information value functions Moravčík et al. (2017); Schmid et al. (2023) without the expensive search in the training time.
@@
In our experiments, we show that SePoT can effectively be combined with RNaD Perolat et al. (2022). We show that using search in head-to-head play in imperfect information Goofspiel and Battleships improves the performance over the network trained with RNaD. Furthermore, we demonstrate that the exploitability of policy derived from the search is lower than that of the policy network from RNaD in a small version of imperfect information Goofspiel and comparable in the standard benchmark game Leduc hold’em. Lastly, we show how using history critic to perform a safe search instead of keeping the statistics from previous searches affects the quality of the final solution.
@@
Our main contributions are: (1) Adding search on top of the policy-gradient method without requiring additional training trajectories. (2) Introducing search technique that may be performed anywhere in the game without needing to perform search in previous parts of the game. (3) Establishing a practical connection between the multi-valued states and the value functions. (4) Showing that the additional search improves the performance of RNaD.
2. Motivation
Figure 1. Current state in a smaller version of Battleships and all possible states sharing the same public information
@@ 2. Motivation
Figure 1. Current state in a smaller version of Battleships and all possible states sharing the same public information
@@
Consider a smaller version of Battleships, in which two players take turns secretly placing three ships of sizes 3, 2, and 2 on a square grid of size 5. Players take turns shooting on an opponent’s board after all ships are placed. The player that first hits all tiles occupied by the opponent’s ships wins the game. Let us suppose we are in the middle of the game, and the current boards are shown in Figure 1. Each player knows their ship’s position but does not know the opponent’s. Based on the public shooting actions up until this point, we know that there are only two possible positions for the player and the opponent. The player may have either placed the ship horizontally with the upper left corner at position A1 or vertically with the upper left corner at position A2. Similarly, the opponent may have placed the ship horizontally with the upper left corner at position A1 or vertically with the upper left corner at position A3. Possible positions are shown on each board with the dashed rectangle. There are four possible states, that is, all possible combinations between the positions across both players, as shown in Figure 0(c). The player cannot distinguish between states $w_{1}$ and $w_{2}$, neither between $w_{3},w_{4}$. Similarly, the opponent cannot distinguish between $w_{1},w_{3}$ and between $w_{2},w_{4}$. Shooting anywhere else than A1, A2, A3, B3, and C3 does not make sense for a rational player. However, training RNaD for three days results in a policy that shoots to these rational tiles with a probability slightly above $40\%$.
@@
Employing additional search should improve this policy, but doing so naively in imperfect information games introduces multiple issues. The first issue is that starting search only from states that are possible from the player’s point of view, which are $w_{1}$ and $w_{2}$, would search only in those parts of the game where the player placed the ship at A1-A2. This means that the player incorrectly assumes the opponent knows the ship’s position during this search. The outcome of the search would be that the game always ends in 2 turns because the opponent would rationally shoot at the ship. Since the player requires at least three turns to win, it could never win against this informed opponent, and its strategy would be to play randomly. However, in the real gameplay scenario, the opponent lacks knowledge of the player’s ship’s location, providing the first player with a genuine opportunity to win. Nevertheless, this issue can be easily fixed by starting the search from all four possible states, which does not make the same error.
@@
The second issue is more profound. In practice, for many different reasons, the player may incorrectly assume that the opponent is slightly more likely to place his ship at position A1 instead of C3. In such a case, the player would always shoot at either of tiles A1, A2, or A3 since it provides more reward in expectation from its point of view. This leads to a strategy that could be exploited by an adversarial opponent who would never put the ship there in the first place. This is a well-studied problem in imperfect information games Burch et al. (2014); Moravcik et al. (2016); Brown and Sandholm (2017); Brown et al. (2020). One possible solution during search is to assume that the opponent can play any strategy in earlier parts of the game. This assumption then allows computing a more robust strategy against possible play from the game’s previous and subsequent parts Burch et al. (2014); Moravcik et al. (2016); Brown and Sandholm (2017).
@@
The third issue for applying full search in the situation in Figure 1 is not tractable since there are more than $10^{9}$ possible terminal histories. Using a depth-limited search with a value function at the depth limit enables search even with much less computational resources Kovařík et al. (2023); Moravčík et al. (2017).
3. Background
@@
Factored-observation stochastic game (FOSG) Kovařík et al. (2022) is a tuple $\mathcal{G}=(\mathcal{N},\mathcal{W},p,w^{0},\mathcal{A},\mathcal{T},\mathcal{R},\mathcal{O})$, where $\mathcal{N}=\{1,\dots,N\}$ is a player set, $\mathcal{W}$ is the set of states and $w^{0}$ is an initial state, $p:\mathcal{W}\rightarrow 2^{\mathcal{N}}$ is a player function, which for given state $w\in\mathcal{W}$ assigns currently acting players from $\mathcal{N}$, $\mathcal{A}=\Pi_{i\in\mathcal{N}}\mathcal{A}_{i}$ is the space of joint actions. We denote $\mathcal{A}_{i}(w)$ legal actions for player $i$ in world state $w$. $\mathcal{T}:\mathcal{W}\times\mathcal{A}\rightarrow\mathcal{W}$ is the transition functiion, $\mathcal{R}:\mathcal{W}\times\mathcal{A}\rightarrow\mathbb{R}^{\mathcal{N}}$ is the reward function and $\mathcal{O}:\mathcal{W}\times\mathcal{A}\times\mathcal{W}\rightarrow\mathbb{O}$ is the observation function. $\mathcal{O}$ is factored as $\mathcal{O}=(\mathcal{O}_{0},\mathcal{O}_{1},\dots\mathcal{O}_{N})$, where $\mathcal{O}_{0}$ are a public observations given to all players and $\mathcal{O}_{i}$ are a private observations of player $i$. In this work, we focus on two-player zero-sum games, where $\mathcal{N}=\{1,2,c\}$ and $\mathcal{R}_{1}(w,a)=-\mathcal{R}_{2}(w,a)\;\forall\mathcal{W}\times\mathcal{A}$. $c$ denotes a chance player who has publicly announced a non-changeable policy throughout the game. Moreover, we study only games where neither player forgets information once it receives it. This is a standard assumption by most of the algorithms, and it is called perfect recall.
@@
Trajectory $\tau=w^{0}a^{0}w^{1}a^{1}\dots w^{t}\in(\mathcal{W}\mathcal{A})^{*}\mathcal{W}$ for which $a^{l}\in\mathcal{A}(w^{l})$ and $w^{l+1}=\mathcal{T}(w^{l},a^{l})$ is a finite sequence of states and actions in the game. Trajectory utility for player $i$ is a cumulative reward $u_{i}(\tau)=\sum_{t=0}^{l-1}\mathcal{R}_{i}(w^{t},a^{t})$. We use $\mathcal{H}$ to denote histories, which are trajectories that start in the initial state $w^{0}$. $\mathcal{Z}=\{z\in\mathcal{H}|z\text{ is terminal}\}$ are histories in which neither player has an action to play, effectively ending the game. $h\sqsubseteq h^{\prime}$ means that $h^{\prime}$ extends $h$ and $h^{0}$ is an initial history, that corresponds to the initial state $w^{0}$, before any player played any action. Each history ends with some world state $w^{l}$. Therefore we denote $\mathcal{R}_{i}(h,a)=\mathcal{R}_{i}(w^{l},a)$ as a reward given to player $i$, if in history $h\in\mathcal{H}$ the players play joint action $a$. Since player $i$ does not observe the state of the game directly, but only observes its actions $\mathcal{A}_{i}$, private observations $\mathcal{O}_{i}$ and public observations $\mathcal{O}_{0}$, it cannot distinguish between multiple different histories. We say that these indistinguishable histories are in the same infoset $s_{i}\in\mathcal{S}_{i}$. If two histories $h,h^{\prime}\in\mathcal{H}$ belong to the same infoset $s_{i}$, then player $i$ has to have the same actions in these histories and $s_{i}(h)$ denotes an infoset of player $i$ in history $h$. We will use $\mathcal{A}_{i}(s_{i})$ to denote these available actions in infoset $s_{i}$. Furthermore, we will use a notion of public-states $s_{0}\in\mathcal{S}_{0}$, which may be viewed as an information set for an external player which does not act in the game. Two histories belong to the same public state if they have the same sequence of public observations from the initial state. Each public state is made from multiple information sets for a given player $i$. We will use $\mathcal{S}_{i}(s_{0})$ to denote all infoset that are in public state $s_{0}$ and $\mathcal{H}(s)$ to denote all histories in an infoset $s$.
@@
Expected utility, if all players play according to strategy profile $\pi$ from history $h$ onward is defined as $u^{\pi}_{i}(h)=\mathbb{E}_{\tau\sim\pi|h}u_{i}(\tau)$. $P^{\pi}(h)=\prod_{h^{\prime}aw\sqsubseteq h}\prod_{j\in\mathcal{N}}\pi_{j}(s_{j}(h^{\prime}),a_{j})$ is a probability of reaching history $h$ if all players play according to strategy profile $\pi$. This may be separated into $P^{\pi}(h)=P^{\pi}_{i}(h)P^{\pi}_{-i}(h)$, where $P^{\pi}_{i}(h)$ is only a players $i$ contribution to reaching history $h$, similarly $P^{\pi}_{-i}(h)$ is a contribution of all the others players except $i$, which is called counterfactual reach probability. Counterfactual values for infoset $s_{i}$ is a sum of expected utilities of each history $h$ within the same infoset, weighted by the counterfactual reach probability of this history $v_{i}^{\pi}(s_{i})=\frac{\sum_{s_{i}(h)=s_{i}}P^{\pi}_{-i}(h)u_{i}^{\pi}(h)}{\sum_{s_{i}(h)=s_{i}}P^{\pi}_{-i}(h)}$.
@@
Best response against policy $\pi_{i}$ of a player $i$ is a policy $\pi_{-i}^{{BR}}\in{BR}_{i}(\pi_{i})$, where $u^{(\pi_{i},\pi^{{BR}}_{-i})}_{i}(h^{0})\geq u^{(\pi_{i},\pi_{-i}^{\prime})}_{i}(h^{0})$ for all $\pi_{-i}^{\prime}$. If all players play a best response to each other, then the resulting policy profile $\pi^{{NE}}$ is called Nash equilibrium. The usual metric to compare strategies in two-player zero-sum games is an exploitability $\mathcal{E}(\pi_{i})=u^{(\pi_{i},\pi^{BR}_{-i})}_{-i}(h^{0})-u^{\pi^{NE}}_{-i}(h^{0})$, that is how much utility can opponent get if it plays best response to the player’s strategy, compared to the utility of Nash equilibrium. If player $i$ plays Nash equilibrium, then the exploitability is 0.
3.1. Depth-limited solving

### 06 Feb 2024 v2

@@
Abstract
Search in test time is often used to improve the performance of reinforcement learning algorithms. Performing theoretically sound search in fully adversarial two-player games with imperfect information is notoriously difficult and requires a complicated training process. We present a method for adding test-time search to an arbitrary policy-gradient algorithm that learns from sampled trajectories. Besides the policy network, the algorithm trains an additional critic network, which estimates the expected values of players following various transformations of the policies given by the policy network. These values are then used for depth-limited search. We show how the values from this critic can create a value function for imperfect information games. Moreover, they can be used to compute the summary statistics necessary to start the search from an arbitrary decision point in the game. The presented algorithm is scalable to very large games since it does not require any search during train time. We evaluate the algorithm’s performance when trained along Regularized Nash Dynamics, and we evaluate the benefit of using the search in the standard benchmark game of Leduc hold’em, multiple variants of imperfect information Goofspiel, and Battleships.
1 Introduction
@@
Despite their success, the neural networks can occasionally lead the agent to make suboptimal decisions that an adversarial opponent may exploit. Wang et al. (2023) have illustrated how policy networks trained via AlphaZero-style algorithms in Go are susceptible to exploitation, as evidenced by winning against the trained policy network 97% of the time. Similarly, Perolat et al. (2022) report that their trained agents display evident gameplay errors, necessitating the application of multiple heuristics to rectify these errors. However, there are principled ways of using search to avoid these mistakes, as evidenced by AlphaZero (Silver et al., 2018), DeepStack (Moravčík et al., 2017), and Student of Games (Schmid et al., 2023).
@@
Algorithms that use policy-gradient methods in imperfect information games like Regularized Nash Dynamics (RNaD) (Perolat et al., 2022, 2021) rely on directly applying the neural network’s policy during gameplay. Adding search to these methods could improve the strategy and mitigate some mistakes. However, this enhancement comes at the expense of turning model-free algorithms into model-dependent ones during gameplay.
@@
Adding search to policy-gradient methods in imperfect information games presents multiple challenges that are not present in a perfect information case, which makes algorithms like AlphaZero unusable in this setting. First, the imperfect information search must be performed not only from the current unobserved state of the game, but from all possible states that share the same public information amongst all players. Second, the state values depend on beliefs players have about the current information of their opponents. Thirdly, the opponent’s strategy is unknown, and the optimal strategies are not unique. Hence, we need to optimize over all possible strategies the opponent may deploy both in the remainder of the game as well as unobserved past actions.
@@
In this paper, we introduce the algorithm Multi-agent Search with Policy Transformations (SePoT), which builds upon the concept of multi-valued states (Brown et al., 2018). SePoT trains a policy transformation critic network that estimates the expected value of policies derived from the policy network. This critic is trained alongside a policy network with a policy-gradient technique without requiring any additional trajectories beyond those sampled by the original policy-gradient algorithm. The critic network is trained with an off-policy V-trace that estimates expected values against a multitude of potential opponent strategies derived from the policy represented by the policy network. We show that this critic provides sufficient information to perform a safe search (Šustr et al., 2021) in imperfect information games.
@@
Furthermore, our approach enables safe search at any point within a game, utilizing the information provided by the introduced critic network. This eliminates the need for conducting search in all decision points from the root until the current decision point, which is often intractable in games with little public knowledge. It is in strict contrast to prior methods (Brown et al., 2018; Moravčík et al., 2017; Schmid et al., 2023). This flexibility allows us to perform search only in game segments sufficiently simple for available compute power, while employing the neural network policy in other parts of the game.
@@
Finally, we establish a connection between value functions and multi-valued states within imperfect information games that enables training previously used imperfect information value functions (Moravčík et al., 2017; Schmid et al., 2023) without the expensive search in the training time.
@@
In our experiments, we evaluate SePoT with RNaD (Perolat et al., 2022). We show that using search in head-to-head play in imperfect-information Goofspiel and Battleships improves the performance over the network trained with RNaD. Furthermore, we demonstrate that the exploitability of policy derived from the search is the same or lower than that of the policy network from RNaD in a small version of imperfect information Goofspiel and the standard benchmark game Leduc hold’em.
@@
Our main contributions are: (1) Adding safe search on top of the policy-gradient method without requiring training time search or additional training trajectories. (2) Introducing search technique that may be performed anywhere in the game without performing search in previous parts of the game. (3) Establishing a practical connection between the multi-valued states and the value functions. (4) Showing that the additional search improves the performance of RNaD.
2 Issues with imperfect information search
@@
Figure 1: Current state in a smaller version of Battleships and all possible states sharing the same public information
@@
Consider a smaller version of Battleships, in which two players have three ships of sizes 3, 2, and 2 on a square grid of size 5. Players take turns shooting on an opponent’s board after all ships are secretly placed. The player that first hits all tiles occupied by the opponent’s ships wins. Suppose we are in the middle of the game, and the current boards are shown in Figure 1. Each player knows their remaining ship’s position, but does not know the opponent’s. Based on the public shooting actions up until this point, we know that there are only two possible positions for the player and the opponent. The player may have either placed the ship at A1-A2 or A2-B2. Similarly, the opponent may have placed the ship at A1-A3 or A3-C3. Possible positions are shown on each board with the dashed rectangle. There are four possible states, that is, all possible combinations between the positions across both players, as shown in Figure 0(c). The player cannot distinguish between states $w_{1}$ and $w_{2}$, neither between $w_{3},w_{4}$. Similarly, the opponent cannot distinguish between $w_{1},w_{3}$ and between $w_{2},w_{4}$. Shooting anywhere else than A1, A2, A3, B3, and C3 does not make sense for a rational player. However, training RNaD for three days results in a policy that shoots to these rational tiles with a probability less than $45\%$.
@@
Employing additional search should improve this policy, but doing so naively in imperfect information games introduces multiple issues. The first issue is that starting search only from states that are possible from the player’s point of view, which are $w_{1}$ and $w_{2}$, would search only in those parts of the game where the player placed the ship at A1-A2. This means that the player incorrectly assumes the opponent knows the ship’s position during this search. The outcome of the search would be that the game always ends in 2 turns because the opponent would shoot at the ship. Since the player requires at least three turns to win, it could never win against this informed opponent, and its strategy would be to play randomly. However, in the real gameplay scenario, the opponent lacks knowledge of the player’s ship’s location, providing the player, who is moving first, with a genuine opportunity to win. Nevertheless, this issue can be fixed by starting the search from all four possible states, which does not make the same error.
@@
The second issue is more profound. In practice, for many different reasons, the player may incorrectly assume that the opponent is slightly more likely to place his ship at position A1 instead of C3. In such a case, the player would always shoot at either tile A1, A2, or A3 since it provides more reward in expectation from its point of view. This leads to a strategy that could be exploited by an adversarial opponent who would never put the ship there in the first place. This is a well-studied problem in imperfect information games. One possible solution during search is to assume that the opponent can play any strategy in earlier parts of the game. This assumption then allows computing a more robust strategy against possible play from the game’s previous and subsequent parts (Burch et al., 2014; Moravcik et al., 2016; Brown and Sandholm, 2017).
@@
The third issue for applying search in the situation in Figure 1 are sizes of required game structures. The full end-game cannot be solved online in reasonable time since it has over $10^{9}$ terminal histories. Also, it is impossible to safely search in earlier decision points, since after the first move there are more than $10^{8}$ possible ship configurations and branching factor of $50$, which menas that even a depth-limited game with look-ahead of one move has more than $10^{10}$ histories. Using a depth-limited search with a value function at the depth limit only later in the game enables search even with much less computational resources (Kovařík et al., 2023).
3 Background
@@
Factored-observation stochastic game (FOSG) (Kovařík et al., 2022) is a tuple $\mathcal{G}=(\mathcal{N},\mathcal{W},p,w^{0},\mathcal{A},\mathcal{T},\mathcal{R},\mathcal{O})$, where $\mathcal{N}=\{1,\dots,N\}$ is a player set, $\mathcal{W}$ is the set of states and $w^{0}$ is an initial state, $p:\mathcal{W}\rightarrow 2^{\mathcal{N}}$ is a player function, which for given state $w\in\mathcal{W}$ assigns currently acting players from $\mathcal{N}$, $\mathcal{A}=\Pi_{i\in\mathcal{N}}\mathcal{A}_{i}$ is the space of joint actions. We denote $\mathcal{A}_{i}(w)$ legal actions for player $i$ in the world state $w$. $\mathcal{T}:\mathcal{W}\times\mathcal{A}\rightarrow\mathcal{W}$ is the transition functiion, $\mathcal{R}:\mathcal{W}\times\mathcal{A}\rightarrow\mathbb{R}^{\mathcal{N}}$ is the reward function and $\mathcal{O}:\mathcal{W}\times\mathcal{A}\times\mathcal{W}\rightarrow\mathbb{O}$ is the observation function. $\mathcal{O}$ is factored as $\mathcal{O}=(\mathcal{O}_{0},\mathcal{O}_{1},\dots\mathcal{O}_{N})$, where $\mathcal{O}_{0}$ are a public observations given to all players and $\mathcal{O}_{i}$ are a private observations of player $i$. In this work, we focus on two-player zero-sum games, where $\mathcal{N}=\{1,2,c\}$ and $\mathcal{R}_{1}(w,a)=-\mathcal{R}_{2}(w,a)\;\forall(w,a)\in\mathcal{W}\times\mathcal{A}$. $c$ denotes a chance player who has publicly announced a non-changeable policy throughout the game. Moreover, we study only games with perfect recall, where neither player forgets any information.
@@
Trajectory $\tau=w^{0}a^{0}w^{1}a^{1}\dots w^{t}\in(\mathcal{W}\mathcal{A})^{*}\mathcal{W}$ for which $a^{l}\in\mathcal{A}(w^{l})$ and $w^{l+1}=\mathcal{T}(w^{l},a^{l})$ is a finite sequence of states and actions in the game. Trajectory utility for player $i$ is a cumulative reward $u_{i}(\tau)=\sum_{t=0}^{l-1}\mathcal{R}_{i}(w^{t},a^{t})$. We use $\mathcal{H}$ to denote histories, which are trajectories that start in the initial state $w^{0}$. $\mathcal{Z}=\{z\in\mathcal{H}|z\text{ is terminal}\}$ are histories in which neither player has an action to play, effectively ending the game. $h\sqsubseteq h^{\prime}$ means that $h^{\prime}$ extends $h$ and $h^{0}$ is an initial history, that corresponds to the initial state $w^{0}$, before any player played any action. Each history ends with some world state $w^{l}$. Therefore, we denote $\mathcal{R}_{i}(h,a)=\mathcal{R}_{i}(w^{l},a)$ as a reward given to player $i$, if in history $h\in\mathcal{H}$ the players play joint action $a$. Since player $i$ does not observe the state of the game directly, but only observes its actions $\mathcal{A}_{i}$, private observations $\mathcal{O}_{i}$ and public observations $\mathcal{O}_{0}$, it cannot distinguish between multiple different histories. We say that these indistinguishable histories are in the same infoset $s_{i}\in\mathcal{S}_{i}$. If two histories $h,h^{\prime}\in\mathcal{H}$ belong to the same infoset $s_{i}$, then player $i$ has to have the same actions in these histories and $s_{i}(h)$ denotes an infoset of player $i$ in history $h$. We will use $\mathcal{A}_{i}(s_{i})$ to denote these available actions in infoset $s_{i}$. Furthermore, we will use a notion of public-states $s_{0}\in\mathcal{S}_{0}$, which may be viewed as an information set for an external player which does not act in the game. Two histories belong to the same public state if they have the same sequence of public observations from the initial state. Each public state is made from multiple information sets for a given player $i$. We will use $\mathcal{S}_{i}(s_{0})$ to denote all infosets that are in public state $s_{0}$ and $\mathcal{H}(s)$ to denote all histories in an infoset $s$.
@@
Expected utility, if all players play according to strategy profile $\pi$ from history $h$ onward is defined as $u^{\pi}_{i}(h)=\mathbb{E}_{\tau\sim\pi|h}u_{i}(\tau)$. $P^{\pi}(h)=\prod_{h^{\prime}aw\sqsubseteq h}\prod_{j\in\mathcal{N}}\pi_{j}(s_{j}(h^{\prime}),a_{j})$ is a probability of reaching history $h$ if all players play according to strategy profile $\pi$. This may be separated into $P^{\pi}(h)=P^{\pi}_{i}(h)P^{\pi}_{-i}(h)$, where $P^{\pi}_{i}(h)$ is only a players $i$ contribution to reaching history $h$, similarly $P^{\pi}_{-i}(h)$ is a contribution of all the other players except $i$, which is called counterfactual reach probability. Counterfactual values for infoset $s_{i}$ is a sum of expected utilities of each history $h$ within the same infoset, weighted by the counterfactual reach probability of this history $v_{i}^{\pi}(s_{i})=\frac{\sum_{s_{i}(h)=s_{i}}P^{\pi}_{-i}(h)u_{i}^{\pi}(h)}{\sum_{s_{i}(h)=s_{i}}P^{\pi}_{-i}(h)}$.
@@
Best response against policy $\pi_{i}$ of a player $i$ is a policy $\pi_{-i}^{{BR}}\in{BR}_{i}(\pi_{i})$, where $u^{(\pi_{i},\pi^{{BR}}_{-i})}_{i}(h^{0})\geq u^{(\pi_{i},\pi_{-i}^{\prime})}_{i}(h^{0})$ for all $\pi_{-i}^{\prime}$. If all players play a best response to each other, then the resulting policy profile $\pi^{{NE}}$ is called Nash equilibrium. The usual metric to compare strategies in two-player zero-sum games is an exploitability $\mathcal{E}(\pi_{i})=u^{(\pi_{i},\pi^{BR}_{-i})}_{-i}(h^{0})-u^{\pi^{NE}}_{-i}(h^{0})$, that is how much utility can opponent get if it plays best response to the player’s strategy, compared to the utility of Nash equilibrium. From definition, if player $i$ plays Nash equilibrium, then the exploitability is 0.
3.1 Depth-limited solving